# Marginal and conditional summary measures: transportability and compatibility

### v1

@@ Abstract
Marginal and conditional summary measures do not generally coincide, have different interpretations and correspond to different decision questions. While these aspects have primarily been recognized for non-collapsible summary measures, they are also problematic for some collapsible measures in the presence of effect modification. We clarify the interpretation and properties of different marginal and conditional summary measures, considering different types of outcomes and hypothetical outcome-generating mechanisms. We describe implications of the choice of summary measure for transportability, highlighting that covariates not conventionally described as effect modifiers can modify population-level treatment effects. Finally, we illustrate existing summary measure incompatibility issues in the context of evidence synthesis, using the case of covariate adjustment methods for indirect treatment comparisons. Because marginal and conditional summary measures do not generally coincide, their naïve pooling in evidence synthesis can produce bias. Almost invariably, care is needed to ensure that evidence synthesis methods are combining compatible summary measures, and this may be easier to ensure with full access to individual patient data.
@@
estimands; evidence synthesis; summary measure; meta-analysis; transportability; indirect treatment comparisons
@@ 1 Introduction
Randomized controlled trials (RCTs) are the “gold standard” research design for estimating causal treatment effects with high internal validity. Randomization ensures that treatment assignment is independent of all measured and unmeasured baseline covariates, thereby guaranteeing that these are balanced in expectation over the trial’s treatment arms. Nevertheless, high internal validity does not necessarily guarantee external validity with respect to a target population of substantive interest. External validity is threatened by differences in the distribution of effect modifiers – covariates inducing variation in treatment effects – between the RCT and the target population (Westreich et al., 2019).
@@
Differences in the distribution of effect modifiers can compromise the transportability or constancy of treatment effects across populations (stat03728, 2014; stat05152, 2015). While said differences are recognized as a fundamental component of treatment effect heterogeneity, a lesser-known source of heterogeneity is the summary (effect) measure used to compare outcomes between treatment arms. Summary measures can be marginal (unconditional) or conditional; directly collapsible (mean and risk difference), collapsible but not directly collapsible (risk ratio) or non-collapsible (odds ratio and hazard ratio) (stat05130, 2015). Definitions of these terms are provided in Table 1.
@@
Marginal and conditional summary measures do not generally coincide and correspond to different decision questions, with discussion over which is more suitable for healthcare decision-making (Phillippo et al., 2021; Remiro-Azócar, 2022b; Phillippo et al., 2025). Whether a summary measure is marginal or conditional has implications on its transportability across populations. In evidence synthesis, the naïve pooling of marginal and conditional measures can produce bias (Remiro-Azócar et al., 2021b, 2022), particularly for analyses using aggregate-level data, as trials may report different summary measures in publications. This issue of “estimand incompatibility” has primarily been recognized for non-collapsible summary measures, where marginal and conditional measures do not coincide, even in the absence of effect modification by the covariates (Daniel et al., 2021). Nevertheless, the issue is also problematic for some collapsible measures – those that are not directly collapsible – in the presence of effect modification (Remiro-Azócar, 2024).
@@
In this article, we will clarify the interpretation of different marginal and conditional summary measures. We do not intend to express a preference towards one or another, but to set out the relevant terminology and illustrate the properties of relevant summary measures. We consider different types of outcomes and outcome-generating mechanisms, with and without effect modification. We describe implications of the choice of summary measure for transportability, highlighting that covariates not conventionally described as effect modifiers can modify population-level treatment effects. We conclude by illustrating existing summary measure incompatibility issues in the context of evidence synthesis, using the case of covariate adjustment methods for indirect treatment comparisons. Almost invariably, care is needed to ensure that evidence synthesis methods are combining compatible summary measures.
@@ 2 Marginal and conditional estimands
An estimand, as defined by the International Council of Harmonisation (ICH) E9 (R1) Addendum, is a precise definition of the treatment effect that is targeted by a clinical trial (Keene et al., 2023). The notion of an estimand as a target of estimation is not new, but its application to pose research questions in registrational RCTs has been encouraged by ICH E9 (R1). The summary measure is only one of the five attributes of estimands, as per ICH E9 (R1). Nevertheless, because our focus is on said attribute, we will use the terms summary measure and estimand interchangeably. Following the potential outcomes framework, these are defined as causal contrasts of any function of the distributions of counterfactual outcomes under different treatments for the same set of subjects (Hernán & Robins, 2020). Because ICH E9 (R1) is aligned with counterfactual language, potential outcomes notation is useful to formally construct the summary measures.
Consider a hypothetical ideal RCT targeting a population of substantive interest. Let $Y^{t}$ denote the potential outcome that might have been observed for an individual assigned to binary treatment $T=t$, with $t\in\{0,1\}$ indicating assignment to the control and the active treatment group, respectively. Each participant in the trial has two potential outcomes, $Y^{1}$ and $Y^{0}$. Assuming no loss to follow up and complete data, one of these is observed in the trial. The other is the outcome hypothetically realized under a different treatment than that actually assigned.
@@
where $g(\cdot)$ is a suitable “link” function, transforming the potential outcome means onto the minus/plus infinity range, and each expectation $E(\cdot)$ is over the distribution of potential outcomes in the trial population. The marginal estimand is typically described as the average treatment effect in the trial population, had everyone in the trial population been assigned active treatment versus control (Austin, 2014). Note that the use of a link function is not strictly necessary for the definition and is usually obviated in the causal inference literature.
@@
While, by virtue of randomization, the marginal estimand within the ideal RCT can be identified from the observed data with minimal assumptions, the identification of conditional estimands within such RCT may require statistical assumptions about correct model specification, particularly with continuous covariates or where there are only a few subjects in given covariate strata (Van Lancker et al., 2022, 2024).
@@
where each expectation $E(\cdot)$ is over the distribution of potential outcomes in the trial population, but conditional on covariate values. The CTEX is the average treatment effect had a subset of individuals, with the same covariate profile, been assigned active treatment versus control. It is often referred to as the conditional average treatment effect (CATE) (Colnet et al., 2023), but we will not use this term to avoid confusion with other average conditional effects introduced later in this section.
To simplify the exposition, we have assumed that there is a single baseline covariate. If this is binary or categorical, the CTEX can be interpreted as a subgroup- or stratum-specific effect. If multiple relevant covariates are conditioned on, it comes closer to an individualized or subject-specific effect. The CTEX is of interest to clinicians and individual patients, particularly if subjects with different covariate values benefit differently from treatment. It is arguably of lesser interest for population-level decision-making, especially if the treatment effect does not change with the covariate values; but can be useful when considering decisions for different subgroups or subpopulations if the treatment effect does change.
@@
with $\bar{X}=E(X)$. This reflects the conditional treatment effect for subjects with covariate values equal to their mean. This summary measure may be of interest if the covariate is continuous. Nevertheless, it is not generally a desirable target for population-level decision-making and likely not meaningful for discrete covariates.
@@
It has been argued that the PACTE can be interpreted as the conditional treatment effect at the (covariate) means (Remiro-Azócar et al., 2021a). Nevertheless, as we shall discuss in Section 3, the PACTE only coincides with the CTEM where the CTEX is constant or a linear function of $X$. The PACTE and the CTEM are fundamentally distinct summary measures. Note that while the CTEX is a function of the covariate(s), the CTEM and the PACTE (and the MTE) compress the treatment effect surface into a single aggregate summary.
@@
$=E_{X}\left(g\left(E\left(Y^{1}\mid X=x\right)\right)\right)-E_{X}\left(g\left(E\left(Y^{0}\mid X=x\right)\right)\right),$
by the law of total expectation. Conversely, the PACTE contrasts the conditional expectations of the potential outcomes on the scale imposed by the link function, then takes the average of the contrasts, with the outer expectation over $X$ outside the link functions. The changing order of operations is irrelevant where $g(\cdot)$ is the identity link, but otherwise matters.
@@ 3 Model-based estimands
The summary measure definitions in Equations 1 to 4 have been formulated on an additive scale. Such definitions are “model-free”; despite the use of function $g(\cdot)$, the estimands do not necessarily rely on modeling assumptions and are not necessarily encoded by the coefficients of a model (Van Lancker et al., 2024; Vansteelandt & Dukes, 2022; Mütze et al., 2025). Nevertheless, in practice, the choices of $g(\cdot)$ and the scale of the summary measure are often “model-based”: influenced by modeling preferences or statistical considerations for different outcome types (Van Lancker et al., 2024; Daniel et al., 2021; Mütze et al., 2025).
@@
Admittedly, investigators are free to derive treatment effects on other scales – e.g., one can use a logit link function to model binary outcomes with a logistic regression, then report the treatment effect as a difference of (marginal) risks, using an identity link for summarization. Nevertheless, as is often the case in practice for RCTs (Wei et al., 2024; Mütze et al., 2025), we shall assume that the scale for the target summary measure is determined by the GLM canonical link function that would be used to model the outcome type. Additionally, we suppose that the conditional outcome expectations follow concrete parameterizations with specific functional forms about the effect of covariates on conditional measures. This implies that the conditional estimands are known and can be expressed in simple terms composed of model parameter(s). In reality, the “true” generative function for the conditional outcome expectation will be more complex, and one is likely to have incomplete knowledge on functional forms for the baseline covariates.
Under the GLM framework, we postulate three hypothetical parametric outcome-generating models to explore the behavior of estimands under different generative mechanisms. These are not exhaustive and have been posited for illustrative purposes. The first outcome-generating mechanism induces treatment effect homogeneity across subjects, such that the CTEX on the linear predictor scale is constant. In this scenario, the MTE, CTEM and PACTE are equivalent for collapsible summary measures, but the MTE does not coincide with any of the conditional quantities for non-collapsible summary measures.
The second outcome-generating mechanism induces treatment effect heterogeneity with the CTEX varying as a linear function of $X$. In this scenario, the MTE, CTEM and PACTE are equivalent for directly collapsible summary measures, but the MTE does not coincide with any of the conditional quantities for summary measures that are not directly collapsible. The third outcome-generating mechanism also induces treatment effect heterogeneity but with the CTEX varying as a non-linear – in this case, quadratic – function of $X$. This highlights a situation where the CTEM and the PACTE are not equivalent, regardless of whether there is collapsibility (or direct collapsibility) or not.
@@ 3.1 Homogeneous CTEX
The outcome model with treatment effect homogeneity, denoted the homogeneous illustrative model, specifies the conditional expectation of potential outcome $Y^{t}$ for treatment $T=t$ given $X$ as:
@@
where $g(\cdot)$ is a suitable link function, $\beta_{0}\in\mathbb{R}$ is an intercept quantifying the “baseline risk” on the linear predictor scale (the conditional outcome expectation on such scale for $X=0$ and $T=0$), and the model coefficients $\beta_{X},\beta_{T}\in\mathbb{R}$ quantify conditional predictor-outcome associations. We place no further statistical assumptions, e.g., normality, on the conditional outcome distribution. Because the outcome model does not include a covariate-treatment product (“interaction”) term but does include a term for the main covariate effect, covariate $X$ is said to be purely prognostic (Ballman, 2015). The baseline risk $\beta_{0}$ is typically a proxy for unobserved prognostic factors.
@@
For a given link function, substituting Equation 5 into Equation 2 yields $\textnormal{CTEX}=\beta_{T}$ on the linear predictor scale. In this case, the CTEX is not actually covariate-specific; it is identical across all individuals or subgroups regardless of their covariate value. As the CTEX does not vary with the level of $X$, the model in Equation 5 imposes treatment effect homogeneity on the linear predictor scale. The conditional treatment effect evaluated at the (covariate) mean $\textnormal{CTEM}=\beta_{T}$ is the same as for every other value of the covariate. Because the PACTE averages the constant CTEX across the subjects in the population, $\textnormal{PACTE}=\beta_{T}$.
@@
With respect to the marginal estimand, consider that $g(\cdot)$ is the identity link, such that the outcome-generating model is linear. On the mean difference scale, by the law of total expectation, the marginal treatment effect estimand $\textnormal{MTE}=\beta_{T}$, and the model-based coefficient $\beta_{T}$ coincides with the model-free definition of the marginal estimand and all conditional estimands (Remiro-Azócar, 2024). With the homogeneous linear generative model, the marginal mean difference does not depend on the distribution of the purely prognostic covariate $X$.
@@
Now, consider that $g(\cdot)=\log(\cdot)$, such that the outcome-generating mechanism is log-linear; for instance, a Poisson or negative binomial model. These models also add the logarithm of person-time, $\log(\tau)$, as an offset to the right-hand side of Equation 5. We shall assume that person-time is fixed across subjects, e.g., corresponding to the trial follow-up period, but we return to this caveat later in this section. On the log rate ratio scale, $\textnormal{MTE}=\beta_{T}$, and the model-based coefficient $\beta_{T}$ coincides with the model-free definition of the marginal estimand and all conditional estimands (Remiro-Azócar, 2024; Vellaisamy & Vijay, 2008). With the homogeneous log-linear generative model (and constant person-time), the marginal (log) rate ratio does not depend on the distribution of the purely prognostic covariate $X$.
@@
Finally, consider that $g(\cdot)=\operatorname{logit}(\cdot)$, such that the outcome-generating model is logistic. Because the (log) odds ratio is non-collapsible (Daniel et al., 2021), the marginal summary measure cannot be expressed as a weighted average of individual- or subgroup-level CTEX, even in the absence of treatment effect heterogeneity (Colnet et al., 2023; Huitfeldt et al., 2019). Assuming $\beta_{X}\neq 0$, the marginal and conditional (log) odds ratios are, almost invariably, not equal, with the marginal estimand closer to the null hypothesis – zero on the log odds ratio scale and one on the odds ratio scale – than all the conditional estimands.
@@
Via Jensen’s inequality and the monotonicity of the logit function, it can be shown that: (1) $\beta_{T}>\textnormal{MTE}$ if $\beta_{T}>0$; and (2) $\beta_{T}<\textnormal{MTE}$ if $\beta_{T}<0$ (Colnet et al., 2023). While the model-based coefficient $\beta_{T}$ can be interpreted as a conditional estimand, it cannot be interpreted as a population-average marginal estimand because it is not equal to the model-free definition of the MTE, despite enforcing treatment effect homogeneity across subjects on the (log) odds ratio scale. Moreover, the MTE on such scale does not only depend on $\beta_{T}$, but depends on the full distribution of the purely prognostic covariate $X$, not just its mean (Daniel et al., 2021), as well as the distribution of baseline risk because prognostic factors may not even be known or measured (Phillippo et al., 2025).
@@
Mean differences, risk differences and risk ratios are collapsible. One would think that (log) rate ratios inherit collapsibility from (log) risk ratios, as rates are risks per unit time. However, (log) rate ratios are non-collapsible (Sjölander et al., 2016), unless person-time is not influenced by the outcomes (treatment or baseline covariates), for instance, in a study of non-fatal recurrent events where all subjects remain at risk throughout the trial follow-up period. We will assume that this is the case for ease of exposition, something which allows us to describe (log) rate ratios as (log) risk ratios when discussing count outcomes and log-linear outcome models.
@@
Table 2 presents a taxonomy of marginal estimands based on the homogeneous illustrative models. Figure 1 indicates whether different estimands match for the homogeneous illustrative models, across link functions and summary measures. The CTEX is not displayed because, outside the homogeneous illustrative model, it is a function of the covariates and not a “one number” summary. The shaded squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent by definition. In Table 2 and Figure 1, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
@@
Figure 1: Matrices indicating whether different estimands are equivalent for the homogeneous illustrative models. The shaded squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
Outside the GLM framework, a model and summary measure that are widely used for the analysis of time-to-event outcomes in RCTs are the Cox proportional hazards model and the (log) hazard ratio. While the Cox model uses a logarithmic link to map hazards to the linear predictor, (log) hazard ratios are non-collapsible for similar reasons to (log) rate ratios, due to conditioning on past survival (Sjölander et al., 2016).
The marginal (log) hazard ratio depends on the shape of the baseline hazard and on the full joint distribution of purely prognostic covariates that do not “interact” with treatment, even under a generative model enforcing the constancy of the conditional (log) hazard ratio across covariate values (Martinussen & Vansteelandt, 2013; Daniel et al., 2021; Phillippo et al., 2025). Awkwardly, such generative model would result in non-proportional hazards for the marginal summary measure because proportional hazards cannot hold simultaneously on the marginal and conditional scales (Daniel et al., 2021; Phillippo et al., 2025).

### v3

@@ Abstract
Marginal and conditional summary measures do not generally coincide, have different interpretations and correspond to different decision questions. While these aspects have primarily been recognized for non-collapsible summary measures, they are equally problematic for some collapsible measures in the presence of effect modification. We clarify the interpretation and properties of several marginal and conditional summary measures, considering different types of outcomes and hypothetical outcome-generating mechanisms. We describe implications of the choice of summary measure for transportability, highlighting that covariates not conventionally described as effect modifiers can modify population-level treatment effects. Finally, we illustrate existing summary measure incompatibility issues in the context of evidence synthesis, using the case of covariate adjustment methods for indirect treatment comparisons. Because marginal and conditional summary measures do not generally coincide, their naïve pooling in evidence synthesis can produce bias. Almost invariably, care is needed to ensure that evidence synthesis methods are combining compatible summary measures, and this may be easier to accomplish with full access to individual patient data.
@@
estimand; evidence synthesis; summary measure; meta-analysis; transportability; indirect treatment comparison
@@ 1 Introduction
Randomized controlled trials (RCTs) are the “gold standard” research design for estimating causal treatment effects with high internal validity. Randomization ensures that treatment assignment is independent of all measured and unmeasured baseline covariates, thereby guaranteeing that these are balanced in expectation over the trial’s treatment arms. Nevertheless, high internal validity does not necessarily guarantee external validity with respect to a target population of substantive interest. External validity is threatened by differences in the distribution of effect modifiers; that is, covariates inducing differential treatment effects between the RCT and the target population (Westreich et al., 2019).
@@
Differences in the distribution of effect modifiers can compromise the transportability or constancy of treatment effects across populations (stat03728, 2014; stat05152, 2015). While such differences are recognized as a fundamental component of treatment effect heterogeneity, a lesser-known source of heterogeneity is the summary (effect) measure used to compare outcomes between treatment arms. Summary measures can be marginal (unconditional) or conditional; directly collapsible (mean and risk difference), collapsible but not directly collapsible (risk ratio) or non-collapsible (odds ratio and hazard ratio) (stat05130, 2015). Definitions of these terms are provided in Table 1.
@@
Marginal and conditional summary measures do not generally coincide and correspond to different decision questions, with discussion over which is more suitable for healthcare decision-making (Phillippo et al., 2021; Remiro-Azócar, 2022b; Phillippo et al., 2025). Whether a summary measure is marginal or conditional has implications on its transportability across populations. In evidence synthesis, the naïve pooling of marginal and conditional measures can produce bias (Remiro-Azócar et al., 2021b, 2022), particularly for analyses using aggregate-level data, as trials may report different summary measures. This issue of “estimand incompatibility” has primarily been recognized for non-collapsible summary measures, where marginal and conditional measures do not coincide, even in the absence of effect modification by the covariates (Daniel et al., 2021). The issue is equally problematic for some collapsible measures – those that are not directly collapsible – in the presence of effect modification (Remiro-Azócar, 2024).
@@
In this article, we clarify the interpretation of different marginal and conditional summary measures. We do not intend to express a preference towards one or another, but to set out the relevant terminology and illustrate the properties of relevant summary measures. We consider different types of outcomes and outcome-generating mechanisms, with and without effect modification. We describe implications of the choice of summary measure for transportability, highlighting that covariates not conventionally described as effect modifiers can modify population-level treatment effects. We conclude by illustrating existing summary measure incompatibility issues in the context of evidence synthesis, using the case of covariate adjustment methods for indirect treatment comparisons. Care has to be exercised to ensure that evidence synthesis methods combine compatible summary measures.
@@ 2 Marginal and conditional estimands
An estimand, as defined by the International Council of Harmonisation (ICH) E9 (R1) Addendum, is a precise definition of the treatment effect that is targeted by a clinical trial (Keene et al., 2023). The notion of an estimand as a target of estimation is not new, but its application to pose research questions in registrational RCTs has been encouraged by ICH E9 (R1). The summary measure is only one of the five attributes of estimands, as per ICH E9 (R1). Nevertheless, because our focus is on such attribute, we will use the terms summary measure and estimand interchangeably. Following the potential outcomes framework, these are defined as causal contrasts of any function of the distributions of counterfactual outcomes under different treatments for the same set of subjects (Hernán & Robins, 2020). Because ICH E9 (R1) is aligned with counterfactual language, potential outcomes notation is useful to formally construct the summary measures.
Consider a hypothetical ideal RCT targeting a population of substantive interest. Let $Y^{t}$ denote the potential outcome that might have been observed for an individual assigned to binary treatment $T=t$, with $t\in\{0,1\}$ indicating assignment to the control and the active treatment group, respectively. Each participant in the trial has two potential outcomes, $Y^{1}$ and $Y^{0}$. Assuming no loss to follow up and complete data, one of these is observed in the trial. The other is the outcome hypothetically realized under the other treatment.
@@
where $g(\cdot)$ is a suitable “link” function, transforming the potential outcome means onto the “minus/plus infinity range”, and each expectation $E(\cdot)$ is over the distribution of potential outcomes in the trial population. The marginal estimand is typically described as the average treatment effect in the trial population, had everyone in the trial population been assigned active treatment versus control (Austin, 2014). Note that the use of a link function is not strictly necessary for the definition and is usually obviated in the causal inference literature.
@@
While, by virtue of randomization, the marginal estimand within the ideal RCT can be identified from the observed data with minimal assumptions, the identification of conditional estimands within such RCT may require statistical assumptions about model validity, particularly with continuous covariates or where there are only a few subjects in some covariate strata (Van Lancker et al., 2022, 2024).
@@
where each expectation $E(\cdot)$ is over the distribution of potential outcomes in the trial population, but conditional on covariate values. The CTEX is the average treatment effect had a subset of individuals, with covariate profile $X=x$, been assigned active treatment versus control. It is often referred to as the conditional average treatment effect (CATE) (Colnet et al., 2023), but we will not use this term to avoid confusion with other average conditional effects introduced later in this section.
To simplify the exposition, we assume that there is a single baseline covariate. If this is binary or categorical, the CTEX can be interpreted as a subgroup- or stratum-specific effect. If multiple relevant covariates are conditioned on, it comes closer to an individualized or subject-specific effect. The CTEX is of interest to clinicians and individual patients, particularly if subjects with different covariate values benefit differently from treatment. It is arguably of lesser interest for population-level decision-making, especially if the treatment effect does not change with the covariate values, but can be useful when considering decisions for different subgroups or subpopulations if the treatment effect does vary.
@@
with $\bar{X}=E(X)$. This reflects the conditional treatment effect for subjects with covariate values equal to their mean. While choice of the mean is somewhat arbitrary, as opposed to the median or any other statistic, this summary measure may be of interest if the covariate is continuous. Nevertheless, it is sensitive to the scale of the covariates, not generally a desirable target for population-level decision-making and likely not meaningful for discrete covariates.
@@
It has been argued that the PACTE can be interpreted as the conditional treatment effect at the (covariate) means (Remiro-Azócar et al., 2021a). Nevertheless, as we shall discuss in Section 3, the PACTE coincides with the CTEM only when the CTEX is a linear function of $X$ (or constant). The PACTE and the CTEM are fundamentally distinct summary measures. Note that while the CTEX is a function of the covariate(s), the CTEM and the PACTE (and the MTE) compress the treatment effect surface into a single aggregate summary.
@@
$=g\left(E_{X}\left(E\left(Y^{1}\mid X=x\right)\right)\right)-g\left(E_{X}\left(E\left(Y^{0}\mid X=x\right)\right)\right),$
by the law of total expectation. Conversely, the PACTE contrasts the conditional expectations of the potential outcomes on the scale imposed by the link function, then takes the average of the contrasts, with the outer expectation over $X$ outside the link functions. The changing order of operations is irrelevant when $g(\cdot)$ is the identity link, but otherwise matters.
@@ 3 Model-based estimands
The summary measure definitions in Equations 1 to 4 are formulated on an additive scale. Such definitions are “model-free”; despite the use of function $g(\cdot)$, the estimands do not necessarily rely on modeling assumptions and are not necessarily encoded by the coefficients of a model (Van Lancker et al., 2024; Vansteelandt & Dukes, 2022; Mütze et al., 2025). Nevertheless, in practice, the choices of $g(\cdot)$ and the scale of the summary measure are often “model-based”: influenced by modeling preferences or statistical considerations for different outcome types (Van Lancker et al., 2024; Daniel et al., 2021; Mütze et al., 2025).
@@
Admittedly, the choice of scale for the comparison is up to the analyst, informed by models as well as value judgments from substantive experts. One can use a logit link function to model binary outcomes with a logistic regression, then report the treatment effect as a difference of (marginal) risks, using an identity link for summarization. Nevertheless, as is often the case in practice for RCTs (Wei et al., 2024; Mütze et al., 2025), we shall assume that the scale for the target summary measure is determined by the GLM canonical link function that would be used to model the outcome type. Additionally, we suppose that the conditional outcome expectations follow particular parameterizations with specific functional forms about the effect of covariates on conditional measures. This implies that the conditional estimands are known and can be expressed in simple terms composed of model parameter(s). In reality, the “true” generative function for the conditional outcome expectation will be more complex, and one is likely to have incomplete knowledge on functional forms for the baseline covariates.
Under the GLM framework, we postulate three hypothetical parametric outcome-generating models to explore the behavior of estimands under different generative mechanisms. These are not exhaustive and are posited for illustrative purposes. The first outcome-generating mechanism induces treatment effect homogeneity across subjects, such that the CTEX on the linear predictor scale is constant. In this scenario, the MTE, CTEM and PACTE are equivalent for collapsible summary measures, but the MTE does not coincide with any of the conditional quantities for non-collapsible summary measures.
The second outcome-generating mechanism induces treatment effect heterogeneity with the CTEX varying as a linear function of $X$. In this scenario, the MTE, CTEM and PACTE are equivalent for directly collapsible summary measures, but the MTE does not coincide with any of the conditional quantities for summary measures that are not directly collapsible. The third outcome-generating mechanism also induces treatment effect heterogeneity but with the CTEX varying as a non-linear – assumed quadratic – function of $X$. In this case, the CTEM and the PACTE are not equivalent, regardless of whether there is collapsibility (or direct collapsibility) or not.
@@ 3.1 Homogeneous CTEX
The outcome model with treatment effect homogeneity, called the homogeneous illustrative model, specifies the conditional expectation of potential outcome $Y^{t}$ for treatment $T=t$ given $X$ as:
@@
where $g(\cdot)$ is a suitable link function, $\beta_{0}\in\mathbb{R}$ is an intercept quantifying the “baseline risk” on the linear predictor scale (the conditional outcome expectation on such scale for $X=0$ and $T=0$), and the model coefficients $\beta_{X},\beta_{T}\in\mathbb{R}$ quantify conditional predictor-outcome associations. We place no further statistical assumptions, e.g., normality, on the conditional outcome distribution. Because the outcome model does not include a covariate-treatment product (“interaction”) term but does include a term for the main covariate effect, covariate $X$ is said to be purely prognostic (Ballman, 2015). The baseline risk $\beta_{0}$ is often a proxy for unobserved prognostic factors (assuming that $X=0$ and $T=0$ correspond to a suitable reference value).
@@
For any link function (the link plays no role in the calculus), substituting Equation 5 into Equation 2 yields $\textnormal{CTEX}=\beta_{T}$ on the linear predictor scale. In this case, the CTEX is not actually covariate-specific; the same CTEX applies to every individual and conceivable subgroup, and we have treatment effect homogeneity. The conditional treatment effect evaluated at the (covariate) mean $\textnormal{CTEM}=\beta_{T}$ is the same as for every other value of the covariate. Because the PACTE averages the constant CTEX across the subjects in the population, $\textnormal{PACTE}=\beta_{T}$.
@@
With respect to the marginal estimand, suppose $g(\cdot)$ is the identity link, such that the outcome-generating model is linear. On the mean difference scale, by the law of total expectation, the marginal treatment effect estimand $\textnormal{MTE}=\beta_{T}$, and the model-based coefficient $\beta_{T}$ coincides with the model-free definition of the marginal estimand and all conditional estimands (Remiro-Azócar, 2024). With the homogeneous linear generative model, the marginal mean difference does not depend on the distribution of the purely prognostic covariate $X$.
@@
Now, suppose $g(\cdot)=\log(\cdot)$, such that the outcome-generating mechanism is log-linear; for instance, a Poisson or negative binomial model. These models also add the logarithm of person-time, $\log(\tau)$, as an offset to the right-hand side of Equation 5. We shall assume that person-time is fixed across subjects, e.g., corresponding to the trial follow-up period, but we return to this caveat later in this section. On the log rate ratio scale, $\textnormal{MTE}=\beta_{T}$, and the model-based coefficient $\beta_{T}$ coincides with the model-free definition of the marginal estimand and all conditional estimands (Remiro-Azócar, 2024; Vellaisamy & Vijay, 2008). With the homogeneous log-linear generative model (and constant person-time), the marginal (log) rate ratio does not depend on the distribution of the purely prognostic covariate $X$.
@@
Finally, suppose $g(\cdot)=\operatorname{logit}(\cdot)$, such that the outcome-generating model is logistic. Because the (log) odds ratio is non-collapsible (Daniel et al., 2021), the marginal summary measure cannot be expressed as a weighted average of individual- or subgroup-level CTEX, not even in the absence of treatment effect heterogeneity (Colnet et al., 2023; Huitfeldt et al., 2019). When $\beta_{X}\neq 0$, the marginal and conditional (log) odds ratios are, almost invariably, not equal, with the marginal estimand closer to the null hypothesis – zero on the log odds ratio scale and one on the odds ratio scale – than all the conditional estimands.
@@
Via Jensen’s inequality and the monotonicity of the logit function, it can be shown that: (1) $\beta_{T}>\textnormal{MTE}$ if $\beta_{T}>0$; and (2) $\beta_{T}<\textnormal{MTE}$ if $\beta_{T}<0$ (Colnet et al., 2023). While the model-based coefficient $\beta_{T}$ can be interpreted as a conditional estimand, it cannot be interpreted as a population-average marginal estimand because it is not equal to the model-free definition of the MTE, despite enforcing treatment effect homogeneity across subjects on the (log) odds ratio scale. Moreover, the MTE on such scale does not only depend on $\beta_{T}$ nor on the mean of purely prognostic covariate $X$. It depends on the full distribution of $X$, as well as the distribution of baseline risk because prognostic factors may not even be known or measured (Phillippo et al., 2025).
@@
Mean differences, risk differences and risk ratios are collapsible. One would think that (log) rate ratios inherit collapsibility from (log) risk ratios, as rates are risks per unit time. However, (log) rate ratios are non-collapsible (Sjölander et al., 2016), unless person-time is not influenced by the outcomes (treatment or baseline covariates), for instance, in a study of non-fatal recurrent events where all subjects remain at risk throughout the trial follow-up period. We will assume that this is the case for ease of exposition; this allows us to describe (log) rate ratios as (log) risk ratios when discussing count outcomes and log-linear outcome models.
@@
Table 2 presents a taxonomy of marginal estimands based on the homogeneous illustrative models. Figure 1 indicates whether different estimands match for the homogeneous illustrative models, across link functions and summary measures. The CTEX is not displayed because, outside the homogeneous illustrative model, it is a function of the covariates and not a “one number” summary. The blue squares denote matching estimand values; the dots denote the diagonal, where logically the estimands are equivalent by definition. In Table 2 and Figure 1, the summary measures are on the additive linear predictor scale imposed by the link function used to model the corresponding outcome type.
@@
Figure 1: Matrices indicating whether different estimands are equivalent for the homogeneous illustrative models. The blue squares denote matching estimand values; the dots denote the diagonal, where estimands are equivalent by definition.
Outside the GLM framework, a model and summary measure that are widely used for the analysis of time-to-event outcomes in RCTs are the Cox proportional hazards model and the (log) hazard ratio. While the Cox model uses a logarithmic link to map hazards to the linear predictor, (log) hazard ratios are non-collapsible owing to conditioning on past survival, similar to (log) rate ratios (Sjölander et al., 2016). The marginal (log) hazard ratio depends on the shape of the baseline hazard and on the full joint distribution of purely prognostic covariates that do not “interact” with treatment, even under a generative model enforcing the constancy of the conditional (log) hazard ratio across covariate values (Martinussen & Vansteelandt, 2013; Daniel et al., 2021; Phillippo et al., 2025). Such generative model would result in non-proportional hazards for the marginal summary measure because proportional hazards cannot hold simultaneously on the marginal and conditional scales (Daniel et al., 2021; Phillippo et al., 2025).